OpenStudy (billyjean):

AB=34,BC=12, and CD=8. What is the value of x?

OpenStudy (anonymous):

\[\large \frac{12}{8} = \frac{34}{x}\] Cross multiply

OpenStudy (billyjean):

22.6 or 23

OpenStudy (anonymous):

Correct! :) A neater way of writing it would be \[\large 22\frac{2}{3}\]

OpenStudy (billyjean):

The asnwer choices are A)61,B)69,C)154,or D)43

OpenStudy (anonymous):

That's odd. Hmm